2/3(x − 3) = 1/6(8x + 31) + 3 i really connot figure this out for some stupid reason

Answer by rothauserc(4718)

2/3(x − 3) = 1/6(8x + 31) + 3
4/6(x-3) = 1/6(8x+31) + 18/6
multiply both sides of = by 6
4(x-3) = (8x+31) + 18
4x -12 = 8x +49
subtract 4x from both sides of =
-12 = 4x +49
subtract 49 from both sides of =
4x = -61
x = -15.25
check answer
substitute for x in original equation
2/3(-15.25-3) = 1/6(8*(-15.25) +31) + 3
−12.166666667 = −12.166666667
our answer checks